Given f(x)=3x-1 is one to one, find f^-1(x)
Answer by nyc_function(2741)

Let y = f(x)
y = 3x - 1
Switch y with x.
x = 3y - 1
Add 1 to both sides.
x + 1 = 3y
Divide both sides by 3.
(x + 1)/3 = y
Replace y with f^-1(x).
(x + 1)/3 = f^-1(x)